OpenStudy (anonymous):

x^3*x^4 without exponents

OpenStudy (solomonzelman):

\[x^3 \times x^4\] \[a^b \times a^c=a^{(b+c)}\]

OpenStudy (solomonzelman):

\[x^3 \times x^4=x^{3+4}=x^7\]

OpenStudy (solomonzelman):

What do you mean without exponents?

OpenStudy (mertsj):

\[x \times x \times x \times x \times x \times x \times x \]

OpenStudy (unklerhaukus):

\[x^n=\overset{\text{n times}}{x\times x\times\cdots\times x}\]

OpenStudy (solomonzelman):

I see what she meant, okay.